grass cutter inc make a product used to trim lawns the firm has has fixed cost of 100000 per year management expect to sell 2000 units per year and that rate of output total variable cost will be an investment of 200000 if the price is set at 100 what is the target rate of return


1
Expert's answer
2021-06-03T13:04:33-0400

"Soln,"

"FC=100000" P.A

Expected sales"=2000" units P.A

"TVC=200000" P.A

Price"=100"


Target rate of return.

"=" "\frac {FC+Target income}{TVC}"


"=" Expected sales "\times" price"=2000\times100=200000"

"=\frac{100000+200000}{200000}=\frac{300000}{200000}=\frac{3}{2}"


"=1.5"